An ice skater skating along an icy surface starts to coast (stops applying a force to the ice). The 

skater coasts a distance of 95 m in 15.8 s. If the coefficient of kinetic friction between the ice 

and skates was 0.077, what was the initial velocity of the skater?


1
Expert's answer
2021-05-13T09:21:25-0400
"s=v_it-0.5\mu g t^2\\95=(15.8)v_i-0.5(0.077)(9.8)(15.8)^2\\v_i=12\frac{m}{s}"